# On the isotopies of tangles in periodic 3-manifolds using finite covers

### v1

@@ Abstract.
A periodic tangle is a one-dimensional submanifold in $\mathbb{R}^{3}$ that has translational symmetry in one, two or three transverse directions. A periodic tangle can be seen as the universal cover of a link in the solid torus, the thickened torus, or the three-torus, respectively. Our goal is to study equivalence relations of such periodic tangles. Since all finite covers of a link lift to the same periodic tangle, it is necessary to prove that isotopies between different finite covers are preserved. In this paper, we show that if two links have isotopic lifts in a common finite cover, then they are isotopic. To do so, we employ techniques from 3-manifold topology to study the complements of such links.
@@ 1. Introduction
Periodic tangles are complex entanglements of curves embedded in $\mathbb{R}^{3}$ that are periodically repeated in either one, two, or three directions. A periodic tangle can be described as the lift of a link in certain 3-manifolds to its universal cover. More specifically, a singly periodic (SP) tangle can be seen as the lift of a link in the solid torus $S^{1}\times D^{2}$ into ×$\mathbb{R}\times D^{2}\subset\mathbb{R}^{3}$ , as described in [59]. A doubly periodic (DP) tangle can been defined as the lift of a link in the thickened torus $T^{2}\times I$ into $\mathbb{R}^{2}\times I$, where $I$ is a closed interval, as done in [13]. Finally, a triply periodic (TP) tangle can be described as the lift of a link in the 3-torus $T^{3}$ into $\mathbb{R}^{3}$, as presented in [2].
@@
Periodic tangles have been used to model a variety of entangled materials. In chemistry, such systems of entangled chains of filaments or catenanes appear in polymers [48, 49, 50, 47], crystalline materials [46, 15], and the mitochondreal DNA of certain protozoa [57, 30]. Periodic tangles also describe macroscopic materials, such as knitted and woven textiles [20, 21, 43, 40, 17, 16, 37], whose physical properties can often be linked to their topology [36, 58]. In all of these systems, the bulk material closely resembles the universal cover of some a link in the solid torus, thickened torus or three-torus. Instead of dealing directly with periodic tangles in the universal cover $\mathbb{R}^{3}$, many researchers consider more minimal representations using links in 3-manifolds, called motifs, that lift to the periodic tangle in question [20, 21, 43, 40]. Textiles, in particular, are represented by links in the thickened torus. When describing the same universal cover, two different researchers might choose different motifs to study. This can be practical from a scientific point of view, but may prove to be confusing when trying to classify the topology of such tangles. For instance, some topological invariants may evaluate to different values, despite referring to the same system in the universal cover.
@@
The topological equivalence of periodic tangles was first investigated by Grishanov, Meshkov, and Omelchenko in [19], whose prime motivation was the study and classification of doubly periodic textile structures through a diagrammatic theory on an integer lattice  [^1] In particular, one may consider a basis $B$ of $\mathbb{E}^{2}$ and a periodic lattice $\Lambda$ isomorphic to $\mathbb{Z}^{2}$ associated to a doubly periodic tangle diagram $D$. Thus the quotient of $D$ under $\Lambda$ generates a motif diagram of a doubly periodic tangle. Such a motif is not unique as infinitely many different point lattices or different choices of basis for the same lattice may be chosen to quotient $D$.. This approach was generalized to the study of doubly periodic tangles in [13], where global isotopies are discretized on the diagrammatic level as sequences of moves on corresponding motif diagrams. These include three types of isotopes acting on (1) the local level, (2) the choice of basis of the universal cover, and (3) number of periodic units in a fundamental domain.
@@
These latter two notions, which extend for both singly periodic tangles ([59]) and singly periodic tangles ([2]), highlight the main difference between isotopies of classical links embedded in a 3-manifold and those of periodic tangles. Note that all of these approaches consider a diagrammatic theory of periodic tangles, which has led to the construction of topological invariants such as several numerical, polynomial, and finite type invariants [2, 5, 13, 14, 19, 20, 21, 22, 23, 29, 43, 59].
@@
Another interesting approach to understanding the topology of periodic tangles is to study the complement of links embedded in $S^{1}\times D^{2}$, $T^{2}\times I$ and $T^{3}$ to account for periodicity. Given a link $L$ in a 3-manifold $X$, removing a tubular neighborhood of $L$ from $X$ results in a 3-manifold known as the link complement, denoted $X\setminus\!\!\setminus L$. A link complement can be systematically constructed from a diagram. Software such as SnapPy does this by constructing a 3-manifold complement combinatorially, starting with the crossings in a diagram [12]. As a result of the geometrization theorem, such 3-manifold complements can (after cutting along appropriate spheres and tori) given metrics consistent with one of eight uniform geometric structures [61, 51, 53, 52]. It is important to note that most knot and link complements are equipped with hyperbolic structures, which have constant sectional curvature $-1$.
@@
While a knot is completely defined by its complement [18], only the majority of the topology of a link is captured by its complement. (We will address the ambiguity arising here in our discussion of Dehn twists and universal covers of periodic tangles.) The properties of the complement 3-manifold can be used to characterize the topology of the knot or link. These are particularly nice when the knot or link complement is a hyperbolic manifold, as invariants of the hyperbolic complement 3-manifold are also invariants of the knot or link. The hyperbolic volume is a topological invariant of a knot that is not directly captured by knot diagrammatic techniques. For example, interesting examples of hyperbolic 3-manifolds have been generated by links in the thickened torus, as investigated in several studies [1, 9, 10, 32, 33, 34, 35, 63, 64].
@@
As with the diagramatic picture, the three levels of isotopy that act a DP tangle $L$ exist when studying the 3-manifold complement $T^{2}\times I\setminus\!\!\setminus L$. (1) Local isotopies are realized a ambient isotopies of the manifold $T^{2}\times I\setminus\!\!\setminus L$. (2) To change the the basis set of the universal cover, let us consider the structure of $T^{2}\times I$. The manifold $T^{2}\times I$ can be obtained by drilling the tubular neighborhood around a Hopf link in $S^{3}.$ The boundary generated by this surgery are identified with $T^{2}\times\{0\}$ and $T^{2}\times\{1\}$. A Dehn twist in the diagramatic picture acts on the components of the Hopf link. This would change the frame on the $T^{2}\times\{0\}$ and/or $T^{2}\times\{1\}$. However, that does not change the structure of the manifold $T^{2}\times I$. (This is the information we have lost between Dehn twists in the diagramatic picture and the 3-manifold picture.) Thus, we find that isotopies between different basis choices on the universal cover to be trivial in the 3-manifold picture. (3) This paper demonstrates that scale equivalence is an equivalence relation between finite covers. We will also show that this equivalence relation holds for singly periodic and triply periodic tangles.
@@ 1.1. Summary of results
The statement of Theorem 1.1 is surprisingly non-trivial. On one hand, it seems fairly obvious that if two links $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ in $\widetilde{X}$ are isotopic, their images $L_{0}=P(\widetilde{L_{0}})$ and $L_{1}=P(\widetilde{L_{1}})$ should be isotopic. However going the other way introduces several complexities. For example, studies in other three-manifolds, including lens spaces and the thickened Klein bottle, have proved that cover-isotopy does not imply isotopy in general [38, 65].
In the theory of periodic tangles, one of the most challenging problem is the identification of a minimal motif, as discussed in [19] for the case of doubly periodic tangles. A motif is the quotient of a periodic tangle under the generators of its translational symmetries. A minimal motif is a motif that cannot be defined as the finite cover of another motif, up to isotopies and Dehn twists. A link is called split if it has a 2-sphere in its complement separating one or more link components from the others. While this holds for links in our three periodic 3-manifolds, we extend this definition to tangles in the universal cover of these manifolds. If a link $L$ in $S^{1}\times D^{2}$, $T^{2}\times I$, and $T^{3}$ is split, we say the corresponding periodic tangle $L_{\infty}$ is also split. In this case, the corresponding periodic tangle $L_{\infty}$ is also called split. Note that a doubly periodic tangles arising from typical textile structures are non-split. In Section 5, we prove the following theorem:
@@
Let $X$ be $S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Suppose that a periodic tangle $L_{\infty}$ has a motif that is a non-split link in $X$. Then $L_{\infty}$ has a motif $L$ such that for any motif $L^{\prime}$ of $L_{\infty}$, there exists a finite covering map $L^{\prime}\to L$. In particular, $L_{\infty}$ has a minimal motif unique up to homeomorphism.
By Theorem 1.2, it is sufficient to consider isotopy of minimal motifs for the classification of non-split periodic tangles. However, it is not easy in general to confirm the minimality of a motif.
To prove our main theorems, we use the standard decomposition of the complement of a link: the prime decomposition and the JSJ decomposition. Each piece after the decomposition is Seifert fibered or hyperbolic. We will classify the types of JSJ tori and the Seifert fibered pieces, and apply the Mostow rigidity to the hyperbolic pieces. Then we will use the result by Budney [6] for the JSJ decomposition of the complement of a link in $S^{3}$. Because the JSJ decomposition often induces links in $X=S^{1}\times D^{2}$, we shall use these results to prove cover-isotopy for $X=T^{2}\times I$ or $T^{3}$ in full generality. We believe that this novel approach will help develop new invariants of periodic tangles and find applications in other fields including textiles, polymers, molecular chemistry, and k-DNA, amongst others.
@@ 2.1. Links in 3-manifolds, periodic tangles, and minimal motifs
In this paper, we consider links in the following 3-manifolds: $X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$, where $I=[0,1]$ is a closed interval. We are interested in studying the lift of these links to the universal cover of $X$, namely ×$\mathbb{R}\times D^{2}$, $\mathbb{R}^{2}\times I$, or $\mathbb{R}^{3}$. In the universal cover, these links are lifted to tangles with translational periodicity, denoted $L_{\infty}$ and called periodic tangles. Conversely, consider the map $P\colon\widetilde{X}\rightarrow X$ to be a finite covering map. Although $\widetilde{X}$ is homeomorphic to $X$, we distinguish $\widetilde{X}$ from $X$ to avoid confusion. This map can be extended to define finite covers of links in $X$. We say that $\widetilde{L}$ is a finite cover of $L$, if there exists a map $P\colon(\widetilde{X},\widetilde{L})\rightarrow(X,L)$, where $P$ is a finite covering map in the usual sense and $\widetilde{L}=P^{-1}(L)$.
@@
([13]) Two periodic tangles are said to be equivalent if they can be obtained from each other by periodic isotopies and orientation-preserving invertible affine transformations of the space carrying along the periodic tangles.
@@
Periodic isotopies are isotopies in the usual sense that preserve periodicity in the universal cover of $X$. Thus it also applies to any finite covering map. Non-periodic local isotopies are not considered in this study.
@@
If two periodic tangles $L_{\infty}$ and $L^{\prime}_{\infty}$ are equivalent and $L$ is a motif of $L_{\infty}$, then we say that $L$ is also a motif of $L^{\prime}_{\infty}$. A consequence of Definition 2.3 is that any finite cover of a motif of $L_{\infty}$ is also a motif of $L_{\infty}$. Thus finite covers of motifs are equivalent under this definition. One might now ask what other types of motif give rise to equivalent periodic tangles. Just as ambient isotopies between links define equivalence, we can also study homeomorphisms of $X\rightarrow X$. This can include trivial diffeomorphisms and homeomorphisms that correspond to Reidemeister movies, but for these periodic manifolds we have an additional class of homeomorphism. In doubly-periodic tangles, it has been shown that Dehn twists are a homeomorphism of $T^{2}\times I\rightarrow T^{2}\times I$ that give rise to equivalent periodic tangles [19, 13]. This can be extend this to Dehn twists in $S^{1}\times D^{2}$ [59] and $T^{3}$ [2].
Next, we introduce the notion of admissible homeomorphism. More precisely, a homeomorphism $h\colon X\to X$ is called admissible if $h$ preserves (1) the orientation for any periodic manifold $X$ and (2) the longitude, up to isotopy, if $X$ is the solid torus $S^{1}\times D^{2}$, or the boundary components (i.e. the front and back) if $X=T^{2}\times I$.
We now have the tools to translate the notion of equivalence between periodic tangles to the language of homeomorphisms of their motifs. Two periodic tangles are equivalent if and only if two motifs of theirs are related by admissible homeomorphisms. Admissible homeomorphisms include the classical equivalence relation of links up to isotopy in $X$ and up to Dehn twists of $X$, as well as the equivalence between two finite covers of the same motif. This leads us to the following proposition:
@@
Let $L_{0,\infty}$ and $L_{1,\infty}$ be two periodic tangles and $L_{0}$ and $L_{1}$ be two of their motifs, respectively, in $X$. Then, $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent if and only if there exist finite covering maps $P_{0},P_{1}\colon X\to X$ such that there is an admissible homeomorphism $(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$.
@@
In [13], the doubly periodic tangle equivalence was investigated by considering motifs along with their supporting lattices. A generalized Reidemeister theorem for doubly periodic tangles translates ambient isotopies of doubly periodic tangles in $\mathbb{R}^{2}\times I$ to the diagrammatic picture of motifs. It proves that two doubly periodic tangles are equivalent if and only if two motif diagrams are related by a finite sequence of moves, which may include (1) parallel shifts of the longitude-meridian pair of the torus, (2) local isotopy moves, (3) Dehn twists (the action of $SL(2,\mathbb{Z})$ on $T^{2}$), or (4) finite covering maps (called scale equivalence).
@@
Figure 1 illustrates examples of the equivalence relations listed in Remark 2.5. On the top left, a doubly periodic (DP) diagram and its quotient under a periodic lattice, that is a link diagram in the (flat) torus. On the top right, two motif diagrams of the same DP diagram related by Reidemeister moves, as well as their corresponding finite covers, such that periodicity is preserved. At the level of the corresponding link in the thickened torus, this equivalence relation between finite covers is one of the main result of this paper. On the bottom, equivalent motifs of the same DP diagram satisfying the following relations: (a) is a double cover of (b), (b) and (c) are related by Dehn twists, (c) is a double cover or (d), and (d) and (e) are related by Dehn twists. These equivalence relations are also considered from a 3-manifold perspective and highlight the open problem of finding a ‘minimal’ motif, which is also addressed in the last section of this paper.
Figure 1. On the top left, a doubly periodic (DP) diagram and a corresponding (flat) motif diagram. On the top right, motifs diagrams and their finite covers related by Reidemeister moves. On the bottom, motifs of the same DP diagrams related by finite covers and Dehn twists.
@@
A motif $L$ of a periodic tangle $L_{\infty}$ is said to be minimal if it is not a non-trivial finite cover of another motif. In other words, $L$ is minimal if it satisfies the following property. Suppose that $L^{\prime}$ is a link in $X$ that is equivalent to $L$, then every finite covering map $(X,L)\rightarrow(X,L^{\prime})$ is a homeomorphism.

### v2

@@ Abstract.
A periodic tangle is a one-dimensional submanifold in $\mathbb{R}^{3}$ that has translational symmetry in one, two, or three transverse directions. A periodic tangle can be seen as the universal cover of a link in the solid torus, the thickened torus, or the three-torus, respectively. Our goal is to study equivalence relations of such periodic tangles. Since all finite covers of a link lift to the same periodic tangle, it is necessary to prove that isotopies between different finite covers are preserved. In this paper, we show that if two links have isotopic lifts in a common finite cover, then they are isotopic. To do so, we employ techniques from 3-manifold topology to study the complements of such links.
@@ 1. Introduction
Periodic tangles are complex entanglements of curves embedded in $\mathbb{R}^{3}$ that are periodically repeated in either one, two, or three directions. A periodic tangle can be described as the lift of a link in certain 3-manifolds to its universal cover. More specifically, a singly periodic (SP) tangle can be seen as the lift of a link in the solid torus $S^{1}\times D^{2}$ into $\mathbb{R}\times D^{2}\subset\mathbb{R}^{3}$, as described in [68]. A doubly periodic (DP) tangle can be defined as the lift of a link in the thickened torus $T^{2}\times I$ into $\mathbb{R}^{2}\times I$, where $I$ is a closed interval, as done in [15]. Finally, a triply periodic (TP) tangle can be described as the lift of a link in the 3-torus $T^{3}$ into $\mathbb{R}^{3}$, as presented in [3].
@@
Periodic tangles have been used to model a variety of entangled materials. In chemistry, such systems of entangled chains of filaments or catenanes appear in polymers [59, 58, 57, 56], crystalline materials [55, 16], and the mitochondrial DNA of certain protozoa [66, 37]. Periodic tangles also describe macroscopic materials, such as knitted and woven textiles [23, 24, 52, 48, 17, 18, 45], whose physical properties can often be linked to their topology [44, 67]. In all of these systems, the bulk material closely resembles the universal cover of some link in the solid torus, the thickened torus, or the 3-torus. Instead of dealing directly with periodic tangles in the universal cover $\mathbb{R}^{3}$, many researchers consider more minimal representations using links in 3-manifolds, called motifs, that lift to the periodic tangle in question [23, 24, 52, 48]. Textiles, in particular, are represented by links in the thickened torus. When describing the same universal cover, two different researchers might choose different motifs to study. This can be practical from a scientific point of view, but it may prove to be confusing when trying to classify the topology of such tangles. For instance, some topological invariants may evaluate to different values, despite referring to the same system in the universal cover.
@@
The topological equivalence of periodic tangles was first investigated by Grishanov, Meshkov, and Omelchenko in [22], whose prime motivation was the study and classification of doubly periodic textile structures through a diagrammatic theory on an integer lattice  [^1] In particular, one may consider a basis $B$ of $\mathbb{E}^{2}$ and a periodic lattice $\Lambda$ isomorphic to $\mathbb{Z}^{2}$ associated to a doubly periodic tangle diagram $D$. Thus, the quotient of $D$ under $\Lambda$ generates a motif diagram of a doubly periodic tangle. Such a motif is not unique, as infinitely many different point lattices or different choices of basis for the same lattice may be chosen to quotient $D$.. This approach was generalized to the study of doubly periodic tangles in [15], where global isotopies are discretized on the diagrammatic level as sequences of moves on the corresponding motif diagrams. These include three types of isotopies:
  1. (1)
    acting on the local level,
  2. (2)
    the choice of a basis of the universal cover, and
  3. (3)
    the number of periodic units in a fundamental domain.
@@
These latter two notions, which extend to both singly periodic tangles [68] and triply periodic tangles [3], highlight the main difference between isotopies of classical links embedded in a 3-manifold and those of periodic tangles. Note that all of these approaches consider a diagrammatic theory of periodic tangles, which has led to the construction of topological invariants such as several numerical, polynomial, and finite type invariants [3, 6, 14, 15, 22, 23, 24, 25, 26, 36, 52, 68].
@@
Another interesting approach to understanding the topology of periodic tangles is to study the complement of links embedded in $S^{1}\times D^{2}$, $T^{2}\times I$, and $T^{3}$ to account for periodicity. Given a link $L$ in a 3-manifold $X$, removing a tubular neighborhood of $L$ from $X$ results in a 3-manifold known as the link complement, denoted $X\setminus\!\!\setminus L$. A link complement can be systematically constructed from a diagram. Software such as SnapPy does this by constructing a 3-manifold complement combinatorially, starting with the crossings in a diagram [13]. As a result of the geometrization theorem, such 3-manifold complements can (after cutting along appropriate spheres and tori) be given metrics consistent with one of eight uniform geometric structures [70, 60, 62, 61]. It is important to note that most knot and link complements are equipped with hyperbolic structures, which have constant sectional curvature 1-1.
@@
While a knot is completely defined by its complement [21], only the majority of the topology of a link is captured by its complement. (We will address the ambiguity arising here in our discussion of Dehn twists and covers of motifs.) The properties of the complement 3-manifold can be used to characterize the topology of the knot or link. These are particularly nice when the knot or link complement is a hyperbolic manifold, as invariants of the hyperbolic complement 3-manifold are also invariants of the knot or link. The hyperbolic volume is a topological invariant of a knot that is not directly captured by knot diagrammatic techniques. For example, two-periodic weft-knitted textiles were investigated using hyperbolic structures in [39]. In addition, interesting examples of hyperbolic 3-manifolds have been generated by links in the thickened torus and the 3-torus, as investigated in several studies [1, 10, 11, 31, 30, 42, 40, 41, 43, 72, 73].
As with the diagrammatic picture, the three levels of isotopy that act on a DP tangle $L$ exist when studying the 3-manifold complement $T^{2}\times I\setminus\!\!\setminus L$.
  1. (1)
    Local isotopies are realized by ambient isotopies of the manifold $T^{2}\times I\setminus\!\!\setminus L$.
  2. (2)
@@
    To change the basis set of the universal cover, let us consider the structure of $T^{2}\times I$. The manifold $T^{2}\times I$ can be obtained by drilling the tubular neighborhood around a Hopf link in $S^{3}.$ The boundary generated by this surgery are identified with $T^{2}\times\{0\}$ and $T^{2}\times\{1\}$. A Dehn twist in the diagrammatic picture acts on the components of the Hopf link. This would change the frame on $T^{2}\times\{0\}$ and/or $T^{2}\times\{1\}$. However, that does not change the structure of the manifold $T^{2}\times I$. (This is the information we have lost between Dehn twists in the diagrammatic picture and the 3-manifold picture.) Thus, we find that isotopies between different basis choices on the universal cover are trivial in the 3-manifold picture.
  3. (3)
    This paper demonstrates that scale equivalence is an equivalence relation between finite covers. We will also show that this equivalence relation holds for singly periodic and triply periodic tangles.
@@ 1.1. Summary of results
[…]
Two braids $\alpha$ and $\beta$ have isotopic closures in the solid torus if and only if $\alpha$ and $\beta$ are conjugate in the braid group $B_{k}$ [35, Theorem 2.1]. Hence Theorem 1.1 implies the following result of Gonzalez-Meneses [20]: If $\alpha$ and $\beta$ in $B_{k}$ satisfy that $\alpha^{n}=\beta^{n}$ for some $n>0$, then $\alpha$ and $\beta$ are conjugate. The proof in [20] relies on the Nielsen-Thurston classification of braids, and so it seems similar to the proof of Theorem 1.1.
In the theory of periodic tangles, one of the most challenging problems is the identification of a minimal motif, as discussed in [22] for the case of doubly periodic tangles. A motif is the quotient of a periodic tangle under the generators of its translational symmetries. A minimal motif is a motif that cannot be a finite cover of another motif. A link is called split if it has a 2-sphere in its complement separating one or more link components from the others. While this is defined for links in our three periodic 3-manifolds, we extend this definition to tangles in the universal cover of these manifolds. If a link $L$ in $S^{1}\times D^{2}$, $T^{2}\times I$, and $T^{3}$ is split, we say that the corresponding periodic tangle $L_{\infty}$ is also split. Note that doubly periodic tangles arising from typical textile structures are not split. In Section 5, we prove the following theorem:
@@
Let $L_{\infty}$ be a non-split periodic tangle with a motif $L\subset X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$. Suppose that each Seifert fibered JSJ piece $M$ of the complement of $L$ such that the homomorphism $\iota_{*}\colon H_{1}(M,\mathbb{Z})\to H_{1}(X,\mathbb{Z})$ induced by the inclusion map is non-trivial can be re-embedded as the complement of $T_{0}(p,0)$, $T_{1}(0,q)$, $T_{2}(p,q)$, or $T_{3}(p,q,r)$. Then $L_{\infty}$ has a motif $L_{\min}$ such that for any motif $L^{\prime}$ of $L_{\infty}$, there exists a finite covering map $L^{\prime}\to L_{\min}$. In particular, $L_{\infty}$ has a minimal motif unique up to homeomorphism.
The notation for Seifert JSJ pieces and re-embedding will be explained in Section 3. The assumption of Theorem 1.2 holds in the case that $L_{\infty}$ is hyperbolic (i.e. it has a motif whose complement admits a hyperbolic structure). By Theorem 1.2, it is sufficient to consider isotopy of minimal motifs for the classification of some types of periodic tangles. However, it is not easy in general to confirm the minimality of a motif. Nonetheless, the symmetry of a hyperbolic periodic tangle can be computed by SnapPy [13], unless it is too complicated.
To prove our main theorems, we use the standard decomposition of the complement of a link: the prime decomposition and the JSJ decomposition. Each piece after the decomposition is Seifert fibered or hyperbolic. We will classify the types of JSJ tori and the Seifert fibered pieces, and apply the Mostow rigidity theorem to the hyperbolic pieces. Then we will use the result of Budney [7] for the JSJ decomposition of the complement of a link in $S^{3}$. Because the JSJ decomposition often induces links in $X=S^{1}\times D^{2}$, we shall use these results to prove cover-isotopy for $X=T^{2}\times I$ or $T^{3}$ in full generality. We believe that this novel approach will help develop new invariants of periodic tangles and find applications in other fields, including textiles, polymers, molecular chemistry, and k-DNA, amongst others.
@@ 2.1. Links in 3-manifolds, periodic tangles, and minimal motifs
In this paper, we consider links in the following 3-manifolds: $X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$, where $I=[0,1]$ is a closed interval. We are interested in studying the lift of these links to the universal cover of $X$, that is, $\mathbb{R}\times D^{2}$, $\mathbb{R}^{2}\times I$, or $\mathbb{R}^{3}$. In the universal cover, these links are lifted to tangles with translational periodicity, denoted $L_{\infty}$ and called periodic tangles. Conversely, consider the map $P\colon\widetilde{X}\rightarrow X$ to be a finite covering map. Although $\widetilde{X}$ is homeomorphic to $X$, we distinguish $\widetilde{X}$ from $X$ to avoid confusion. This map can be extended to define finite covers of links in $X$. We say that $\widetilde{L}$ is a finite cover of $L$, if there exists a map $P\colon(\widetilde{X},\widetilde{L})\rightarrow(X,L)$, where $P$ is a finite covering map in the usual sense and $\widetilde{L}=P^{-1}(L)$.
@@
([15]) Two periodic tangles are said to be equivalent if they can be obtained from each other by periodic isotopies and invertible affine transformations of the space carrying along the periodic tangles that are isotopic to the identity.
@@
Periodic isotopies are isotopies in the usual sense that preserve periodicity in the universal cover of $X$. Thus it also applies to any finite covering map. Non-periodic local isotopies are not considered in this study. Note that periodic lattices are not fixed in this definition. In other words, it is sufficient to preserve a large lattice.
@@
If two periodic tangles $L_{\infty}$ and $L^{\prime}_{\infty}$ are equivalent and $L$ is a motif of $L_{\infty}$, then we say that $L$ is also a motif of $L^{\prime}_{\infty}$. A consequence of Definition 2.3 is that any finite cover of a motif of $L_{\infty}$ is also a motif of $L_{\infty}$. Thus, finite covers of motifs are equivalent under this definition. One might now ask what other types of motifs give rise to equivalent periodic tangles. Just as ambient isotopies between links define equivalence, we can also study homeomorphisms of $X\rightarrow X$. This can include trivial diffeomorphisms and homeomorphisms that correspond to Reidemeister moves, but for these periodic manifolds, we have an additional class of homeomorphisms. In doubly-periodic tangles, it has been shown that Dehn twists are homeomorphisms of $T^{2}\times I\to T^{2}\times I$ that give rise to equivalent periodic tangles [22, 15]. This can be extended to Dehn twists in $S^{1}\times D^{2}$ [68] and $T^{3}$ [3].
Next, we introduce the notion of admissible homeomorphism. More precisely, a homeomorphism $h\colon X\to X$ for $X=S^{1}\times D^{2}$, $T^{2}\times I$, or $T^{3}$ is called admissible if $h$ satisfies the following:
  1. (1)
    The homeomorphism $h$ preserves the orientation of $X$.
  2. (2)
    If $X=S^{1}\times D^{2}$, the homeomorphism $h$ is isotopic to the identity.
  3. (3)
    If $X=T^{2}\times I$, the homeomorphism $h$ preserves the boundary components (i.e. the front and back).
We now have the tools to translate the notion of equivalence between periodic tangles to the language of homeomorphisms of their motifs. Two periodic tangles are equivalent if and only if two motifs of theirs are related by admissible homeomorphisms. If $X=T^{2}\times I$, the admissible homeomorphisms include the classical equivalence relation of links up to isotopy in $X$ and up to Dehn twists of $X$, as well as the equivalence between two finite covers of the same motif. This leads us to the following proposition:
@@
Let $L_{0,\infty}$ and $L_{1,\infty}$ be two periodic tangles and $L_{0}$ and $L_{1}$ be two of their motifs, respectively, in $X$. Then $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent if and only if there exist finite covering maps $P_{0},P_{1}\colon X\to X$ such that there is an admissible homeomorphism $h\colon(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$.
Proof.
Suppose that there exist finite covering maps $P_{0},P_{1}\colon X\to X$ such that there is an admissible homeomorphism $h\colon(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$. Then $h$ lifts to a homeomorphism $h_{\infty}\colon(X_{\infty},L_{0,\infty})\to(X_{\infty},L_{1,\infty})$ that is equivariant under the translational periodicity and isotopic to the identity. Hence $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent.
Conversely, suppose that $L_{0,\infty}$ and $L_{1,\infty}$ are equivalent. Then, by taking composition with an affine transformation, we may assume that there is a homeomorphism $h_{\infty}\colon(X_{\infty},L_{0,\infty})\to(X_{\infty},L_{1,\infty})$ that is equivariant under the translational periodicity of $\mathbb{Z}^{d}$ ($d=1$, 22, or $3$) acting on $X_{\infty}$ and isotopic to the identity. Here $(X,L_{i})$ for each $i=0,1$ is the quotient of $(X_{\infty},L_{i,\infty})$ by a lattice $\Lambda_{i}\subset\mathbb{R}^{d}$. If $\Lambda_{i}\subset\mathbb{Q}^{d}$ for both $i=0,1$, there is a lattice $\widetilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\widetilde{\Lambda}\subset\Lambda_{0}$ and $\widetilde{\Lambda}\subset\Lambda_{1}$. Let $P_{i}\colon X\to X$ denote the finite covering maps that correspond to $\widetilde{\Lambda}\subset\Lambda_{i}$. Then $h_{\infty}$ induces an admissible homeomorphism $h\colon(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$.
Otherwise the closure of the subgroup of $\mathbb{R}^{d}$ generated by $\mathbb{Z}^{d}$ and $\Lambda_{i}$ for $i=0$ or $1$ contains a 1-dimensional subspace $V$. Since $(X_{\infty},L_{i,\infty})$ is invariant under the action of $V$, the periodic tangle $L_{i,\infty}$ consists of parallel straight lines in $X_{\infty}$. In this case, there are finite covering maps $P_{i}\colon X\to X$ such that $P_{0}^{-1}(L_{0})$ and $P_{1}^{-1}(L_{1})$ are isotopic to parallel closed geodesics in $X$ of the same component number. Hence again there is an admissible homeomorphism $h\colon(X,P_{0}^{-1}(L_{0}))\to(X,P_{1}^{-1}(L_{1}))$. ∎
@@
In [15], the doubly periodic tangle equivalence was investigated by considering motifs along with their supporting lattices. A generalized Reidemeister theorem for doubly periodic tangles translates ambient isotopies of doubly periodic tangles in $\mathbb{R}^{2}\times I$ to the diagrammatic picture of motifs. It proves that two doubly periodic tangles are equivalent if and only if two motif diagrams are related by a finite sequence of moves, which includes
  1. (1)
    parallel shifts of the longitude-meridian pair of the torus,
  2. (2)
    local isotopy moves,
  3. (3)
    Dehn twists (the action of $SL(2,\mathbb{Z})$ on $T^{2}$), and
  4. (4)
    finite covering maps (called scale equivalence).
@@
Figure 1 illustrates examples of the equivalence relations listed in Remark 2.5. On the top left, a doubly periodic (DP) diagram and its quotient under a periodic lattice, that is, a link diagram in the (flat) torus. On the top right, two motif diagrams of the same DP diagram related by Reidemeister moves, as well as their corresponding finite covers, such that periodicity is preserved. At the level of the corresponding link in the thickened torus, this equivalence relation between finite covers is one of the main results of this paper. On the bottom, equivalent motifs of the same DP diagram satisfying the following relations:
  • (a) is a double cover of (b),
  • (b) and (c) are related by Dehn twists,
  • (c) is a double cover or (d), and
  • (d) and (e) are related by Dehn twists.
These equivalence relations are also considered from a 3-manifold perspective and highlight the open problem of finding a ‘minimal’ motif, which is also addressed in the last section of this paper.
Figure 1. On the top left, a doubly periodic (DP) diagram and a corresponding (flat) motif diagram. On the top right, motif diagrams and their finite covers are related by Reidemeister moves. On the bottom, motifs of the same DP diagrams related by finite covers and Dehn twists.
@@
A motif $L$ of a periodic tangle $L_{\infty}$ is said to be minimal if it is not a non-trivial finite cover of another motif. In other words, $L$ is minimal if it satisfies the following property. Suppose that $L^{\prime}$ is a link in $X$ that is equivalent to $L$, then every finite covering map $(X,L)\to(X,L^{\prime})$ is a homeomorphism.
In [15], another definition of a minimal motif is given in terms of the associated periodic lattice. In our definition, we have to be aware that a periodic tangle may become more symmetric by a periodic isotopy. The existence and uniqueness of a minimal motif are non-trivial questions, which we will discuss in Section 5. Note that seeking systematic ways to find a minimal motif is still an open problem.